# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v3

@@ Abstract
Diffusion models demonstrate remarkable capabilities in capturing complex data distributions and have achieved compelling results in many generative tasks. While they have recently been extended to dense prediction tasks such as depth estimation and surface normal prediction, their full potential in this area remains under-explored. In dense prediction settings, target signal maps and input images are pixel-wise aligned. This makes conventional noise-to-data generation paradigm inefficient, as input images can serve as more informative prior compared to pure noise. Diffusion bridge models, which support data-to-data generation between two general data distributions, offer a promising alternative, but they typically fail to exploit the rich visual priors embedded in large pretrained foundation models. To address these limitations, we integrate diffusion bridge formulation with structured visual priors and introduce DPBridge, the first latent diffusion bridge framework for dense prediction tasks. Our method presents three key contributions: (1) a tractable reverse transition kernel for diffusion bridge process, enabling maximum likelihood training scheme for better compatibility with pretrained backbones; (2) a distribution-aligned normalization technique to mitigate the discrepancies between the bridge and standard diffusion processes; and (3) an auxiliary image consistency loss to preserve fine-grained details. Experiments across extensive benchmarks validate that our method consistently achieves superior performance, demonstrating its effectiveness and generalization capability under different scenarios.
@@ 1 Introduction
Dense prediction refers to a class of problems in which the goal is to assign a prediction value for every pixel in the input image. It encompasses a variety of fundamental tasks, including depth estimation [56, 39, 15], surface normal prediction [2, 22, 3], semantic segmentation [53, 8, 51], optical flow [49, 23, 54], etc.
@@
Recently, image diffusion models have been applied to dense prediction tasks by framing them as conditional generation problems. Leveraging the ability to capture complex data distributions, diffusion models serve as powerful priors and yield promising results [28, 17, 16]. However, these approaches exhibit inherent limitations stemming from the underlying generative mechanism. Specifically, the denoising process always starts from Gaussian noise, which carries minimal information about the target signal distribution. In such cases, the distributional gap between initial Gaussian prior and target distribution remains large, necessitating more sampling steps and often reliance on manually designed post-hoc guidance [9, 46, 10] or customized generation pipeline [35]. This not only increases computational overhead but can also degrade performance and prolong convergence.
@@
Instead of starting from noise, a more effective alternative is to directly initiate the generative process from the input. Since RGB images are pixel-wise aligned with their corresponding dense signal maps, they serve as informative priors. This motivates the exploration of diffusion bridge models [34, 37, 14, 33], another line of computational frameworks based on stochastic processes with fixed start and end states, designed to model probability flows between two given distributions. We argue that diffusion bridge models are more suitable for dense prediction tasks, as their data-to-data generation paradigm naturally leverage the structural correspondence between inputs and outputs, thereby mitigating inefficiencies inherent in conventional noise-to-data scheme. Despite theoretical feasibility, diffusion bridge-based methods [65, 31, 32] have not yet demonstrated competitive performance compared with standard diffusion counterparts. This performance gap largely stems from the fact that most bridge models are trained from scratch via score matching objectives, making them incompatible with commonly used pretrained diffusion backbone. As a result, they fail to exploit the embedded semantic and spatial prior knowledge, limiting their expressiveness and generalization ability in dense prediction settings.
@@
In this paper, we propose Dense Prediction Bridge (DPBridge), the first latent diffusion bridge model tailored for general dense prediction tasks by integrating the strengths of diffusion bridge modeling with visual priors from pretrained backbones. Specifically, we derive a fully tractable bridge process to connect the latent representations of RGB images and their corresponding dense signal maps, and formulate the generation process as a sequence of transformations that progressively synthesize the target maps from input data. Furthermore, we introduce finetuning strategies to better adapt pretrained diffusion backbones for DPBridge construction. Leveraging the rich visual priors not only enhances perceptual performance and generalization capability, but also significantly accelerates training convergence. Figure 1 provides a high-level overview of DPBridge framework and illustrates its effectiveness across three representative dense prediction tasks: depth estimation, surface normal prediction, and semantic segmentation. These results demonstrate that our framework is both theoretically general and practically adaptable to diverse dense prediction scenarios.
@@
  • We propose DPBridge, a generative framework to address dense prediction tasks by formulating them as instances of latent diffusion bridges.
@@
  • We introduce a distribution-aligned normalization technique that mitigates distributional discrepancies between the bridge and standard diffusion processes, allowing for more effective finetuning. In addition, by incorporating auxiliary image consistency loss, we can better preserve fine-grained spatial details often lost in latent representations, enhancing per-pixel prediction accuracy.
  • Our method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection.
@@ 2 Related Works
Dense Prediction. Dense predictions can be characterized as conditional distribution modeling and common solutions are categorized into discriminative and generative ones. Discriminative methods directly map input data to structured outputs in a feed-forward manner. [39] introduces transformer architecture and fusion mechanism to better capture both local and global context feature. [56] explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness. [2] incorporates per-pixel ray directions and models the relative rotations between neighboring pixels. Recently, diffusion-based generative methods have been applied in this field and showed remarkable performance. [28] finetunes pretrained diffusion backbone, leveraging the rich visual knowledge within to enhance depth estimation. [16] takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions. [19] exploits flow matching to facilitate the distribution shift between images and depth maps. Our work follow conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in iterative denoising manner.
@@
Diffusion and Bridge Models. Diffusion models [20, 47] are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes. They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis [41, 21], controlled editing [62, 4], inverse problem solving [9, 46], 3D assets generation [55, 63], and human-centered applications [26, 25, 24].
@@
Diffusion bridge models have gained increasing attention in generative modeling in recent years [37, 34, 65, 33, 14]. They offer a novel perspective for translating between distributions and enable a more effective data-to-data generation paradigm. Recent works have explored different types of bridge processes and their applications. [32, 31, 60] develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution. [11] proposes a modified inference procedure that enforces data consistency. [7] introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency.
@@
To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness. Our work aims to address this research gap. Conceptually, DPBridge is most related to DDBM [65], but the two differ in several key aspects: DDBM is trained from scratch using a score matching objective, whereas DPBridge is finetuned from large pretrained diffusion backbones using a maximum likelihood objective. In addition, we introduce training strategies tailored for efficient finetuning, enabling better use of visual priors and significantly improving performance on dense prediction settings.
@@ 3.1 Diffusion Process
A diffusion process =$\mathbb{Q}=\{\mathbf{z}_{t}\in\mathbb{R}^{d}:t\in[0,T]\}$ on space $\mathbb{R}^{d}$ follows the stochastic differential equation of form
@@
where $\mathbf{w}_{t}$ is a standard Wiener process, $f:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ and $g:[0,T]\rightarrow\mathbb{R}$ are the drift and diffusion coefficients that describe the evolution of this stochastic process. The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\sigma_{t}$ depends on the design choice of drift and diffusion coefficients. It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$.
@@ 3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that is conditioned to end at a specified terminal point at timestep $T$. One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints. We can use Doob’s h-transform technique [36, 42] to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value. The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
@@
$m_{t}=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right),\quad n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}},\quad\sigma_{t}=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}}$
where $\text{SNR}_{t},\alpha_{t},\sigma_{t}^{2}$ are all ingredients defined in the reference process $\mathbb{Q}$.
@@ 4.1 Problem Formulation
We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ denotes the target dense signal map (e.g. depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image. In the following sections, we’ll elaborate the details of DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
@@ 4.2 Dense Prediction Bridge
Given paired training data, a straightforward approach is to construct the diffusion bridge model from scratch, following the formulation in Section 3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more, it lacks the ability to harness rich visual prior embedded in the foundation models pretrained with large-scale data,which have been shown to significantly enhance both perceptual quality and generalization capability in dense prediction tasks. To address these limitations and fully exploit the visual prior, our DPBridge is build upon pretrained diffusion backbones through finetuning. We explore several strategies to handle the misalignment between bridge process and standard diffusion process, facilitating more efficient knowledge transfer. An overview of our training pipeline is provided in Figure 2, followed by a detailed explanation of each component.
@@ 4.3 Latent Space Implementation
We follow common practice to perform diffusion steps in a compressed yet perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the dense signal map and RGB image are encoded using the VAE encoder, yielding latent representations $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, which serve as the initial and terminal points of the diffusion bridge process. In this way the diffusion bridge process is operated entirely in the latent space. To recover the original signal map, we use the VAE decoder to decode the predicted latent $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. For clarity, all subsequent derivations in this paper are conducted in latent space.
@@ 4.4 Training Objective
Maximum Likelihood Loss. A widely adopted approach for learning the reverse process associated with Equation 2 follows the theoretical framework of [1], where a neural network is trained to approximate the score function [65]. In contrast, we observe that in the diffusion bridge setting, the reverse transition kernel admits a tractable analytical form. This insight enables us to adopt a maximum likelihood training paradigm in place of score matching. Since widely used latent diffusion models like Stable Diffusion [41] are pretrained under a likelihood-based loss function, finetuning with a consistent objective is more compatible, which helps preserve prior knowledge and improves training stability.
@@
[…]
@@
For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$. Based on Equation 3, we further derive the mean value term:
@@
$k_{1}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t},\quad k_{2}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot$ $m_{t-1},\quad k_{3}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
@@
Assuming $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})\sim\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ is also a Gaussian distribution with fixed variance $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)=\tilde{\sigma}_{t}^{2}$ for simplicity of derivation. Following the strategy of DDPM [20], we reparameterize the model $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to predict noise instead of posterior mean. Under this formulation, the ELBO objective is equivalent to:

### v4

@@ Abstract
Diffusion models demonstrate remarkable capabilities in capturing complex data distributions and have achieved compelling results in many generative tasks. While they have recently been extended to dense prediction tasks such as depth estimation and surface normal prediction, their full potential in this area remains underexplored. As target signal maps and input images are pixel-wise aligned, the conventional noise-to-data generation paradigm is inefficient, and input images can serve as a more informative prior compared to pure noise. Diffusion bridge models, which support data-to-data generation between two general data distributions, offer a promising alternative, but they typically fail to exploit the rich visual priors embedded in large pretrained foundation models. To address these limitations, we integrate diffusion bridge formulation with structured visual priors and introduce DPBridge, the first latent diffusion bridge framework for dense prediction tasks. To resolve the incompatibility between diffusion bridge models and pretrained diffusion backbones, we propose (1) a tractable reverse transition kernel for the diffusion bridge process, enabling maximum likelihood training scheme; (2) finetuning strategies including distribution-aligned normalization and image consistency loss. Experiments across extensive benchmarks validate that our method consistently achieves superior performance, demonstrating its effectiveness and generalization capability under different scenarios.
@@ 1 Introduction
Dense prediction refers to a class of problems in which the goal is to assign a prediction value for every pixel in the input image. It encompasses a variety of fundamental tasks, including depth estimation [yang2024depth, ranftl2021vision, eftekhar2021omnidata], surface normal prediction [bae2024rethinking, hu2024metric3d, bae2021estimating], semantic segmentation [weber2021step, cheng2022masked, wang2023images], optical flow [teed2020raft, huang2022flowformer, weinzaepfel2023croco], etc.
@@
Dense prediction has traditionally relied on discriminative models [ranftl2021vision, zheng2021rethinking, bae2024rethinking], which map RGB inputs to dense signal maps directly via feed-forward manners. While these methods offer high inference speed, they usually struggle in ambiguous scenarios-such as occluded regions or dynamic surface normals of glossy objects. This limitation is largely due to the lack of uncertainty modeling. Recently, generative models such as diffusions have been applied to mitigate these drawbacks by framing dense prediction as a conditional generation task. Leveraging the ability of modeling complex data distributions, diffusions serve as robust priors and deliver promising performance as different tasks [ke2024repurposing, garcia2024fine, fu2025geowizard]. Nevertheless, diffusion-based generative approaches suffer from inherent limitations rooted in their core mechanism: their denoising process initiates from Gaussian noise—an initial prior that contains minimal information about the target dense signal distribution. In such cases, the distributional gap between initial Gaussian prior and target distribution remains large, necessitating more sampling steps and often reliance on manually designed post-hoc guidance [chung2022diffusion, song2023pseudoinverse, chung2022improving] or customized generation pipeline [meng2021sdedit]. This not only increases computational overhead but can also degrade performance and prolong convergence.
@@
Instead of starting from noise, a more effective alternative is to directly initiate the generative process from the input. Since RGB images are pixel-wise aligned with their corresponding dense signal maps, they shall act as highly informative priors. This motivates the exploration of diffusion bridge models [liu2022let, pavon2021data, de2021diffusion, liu2023learning], another line of computational frameworks based on stochastic processes with fixed start and end states, designed to model probability flows between two given distributions. We argue that diffusion bridge models are more suitable for dense prediction tasks, as their data-to-data generation paradigm naturally leverages the structural correspondence between inputs and outputs, thereby mitigating inefficiencies inherent in the conventional noise-to-data scheme. Despite theoretical feasibility, diffusion bridge-based methods [zhou2023denoising, li2023bbdm, liu20232] have not yet demonstrated competitive performance compared with standard diffusion counterparts. This performance gap primarily arises because most bridge models are trained from scratch via score matching objectives, making them incompatible with the widely adopted pretrained diffusion backbones. Consequently, these methods fail to leverage the embedded semantic and spatial prior knowledge, constraining their expressive power and generalization capability in dense prediction scenarios.
@@
In this paper, we propose Dense Prediction Bridge (DPBridge), the first latent diffusion bridge model tailored for general dense prediction tasks by integrating the strengths of diffusion bridge modeling with visual priors from pretrained backbones. Specifically, we derive a fully tractable bridge process to connect the latent representations of RGB images and their corresponding dense signal maps, and formulate the generation process as a sequence of transformations that progressively synthesize the target maps from input data. Furthermore, we introduce finetuning strategies to better adapt pretrained diffusion backbones for DPBridge construction. Leveraging rich visual priors not only enhances perceptual performance and generalization ability, but also significantly accelerates training convergence. Figure˜1 provides a high-level overview of the DPBridge framework and illustrates its effectiveness across two representative tasks: depth estimation and surface normal prediction. These results demonstrate that our method is both theoretically feasible and practically adaptable to diverse scenarios.
@@
  • We propose DPBridge, a generative framework to address general dense prediction tasks by formulating them as instances of latent diffusion bridges.
@@
  • We introduce finetuning strategies to improve finetuning performance, including a distribution-aligned normalization technique that mitigates distributional discrepancies between the bridge and standard diffusion processes, and an auxiliary image consistency loss to better preserve fine-grained spatial details.
  • Our method shows competitive performance across different tasks in terms of both quantitative metrics and visual inspection.
@@ 2 Related Works
Dense Prediction. Common solutions to dense predictions are categorized into discriminative and generative ones. Discriminative methods directly map input data to structured outputs in a feed-forward manner. [ranftl2021vision] introduces transformer architecture and fusion mechanism to better capture both local and global context features. [bae2024rethinking] incorporates per-pixel ray directions and models the relative rotations between neighboring pixels. Recently, diffusion-based generative methods have been applied in this field and have shown remarkable performance. [ke2024repurposing] finetunes pretrained diffusion backbone, leveraging the rich visual knowledge within to enhance depth estimation. [fu2025geowizard] takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions. [gui2024depthfm] exploits flow matching to facilitate the distribution shift between images and depth maps. Our work follows the conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in an iterative denoising manner.
@@
Diffusion and Bridge Models. Diffusion models [ho2020denoising, song2020score] are a family of generative models that excel at capturing complex data distributions through a sequence of forward and reverse processes. They have driven the advancement of modern generative systems, and have consistently achieved state-of-the-art performance across various domains, such as image and video synthesis [rombach2022high, ho2022video], controlled editing [zhang2023adding, batzolis2021conditional], inverse problem solving [chung2022diffusion, song2023pseudoinverse], 3D assets generation [xiang2024structured, zhang2024clay], and human-centered applications [ji2024jade, ji2024unsupervised, ji20243d].
@@
In recent years, diffusion bridge models have gained growing interest in generative modeling [pavon2021data, liu2022let, zhou2023denoising, liu2023learning, de2021diffusion]. They offer a novel perspective for translating between distributions and enable a more effective data-to-data generation paradigm. Recent works have explored different types of bridge processes and their applications. [liu20232, li2023bbdm, yue2023image] develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution. [chung2024direct] proposes a modified inference procedure that enforces data consistency. [chen2023schrodinger] introduces bridge models to text-to-speech systems, enhancing synthesis quality.
@@
To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness. Our work aims to fill this research gap. Conceptually, DPBridge is most related to DDBM [zhou2023denoising], but the two differ in several key aspects: DDBM is trained from scratch using a score matching objective, whereas DPBridge is finetuned from large pretrained diffusion backbones using a maximum likelihood objective. In addition, we introduce strategies for efficient finetuning, enabling better use of visual priors and significantly improving performance on dense prediction settings.
@@ 3.1 Diffusion Process
A diffusion process $\mathbb{Q}=\{\mathbf{z}_{t}\in\mathbb{R}^{d}:t\in[0,T]\}$ on space $\mathbb{R}^{d}$ follows the stochastic differential equation of form
@@
where $\mathbf{w}_{t}$ is a standard Wiener process, $f:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ and $g:[0,T]\rightarrow\mathbb{R}$ are the drift and diffusion coefficients that describe the evolution of this stochastic process. The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\eta_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\eta_{t}$ depends on the design choice of drift and diffusion coefficients. It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\eta_{t}^{2}$.
@@ 3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that is conditioned to end at a specified terminal point at timestep $T$. Its dual fixed-endpoint characteristics enable a bidirectional distribution transport. Its construction can be derived from a reference diffusion process conditioned on terminal constraints. We can use Doob’s h-transform technique [oksendal2013stochastic, sarkka2019applied] to modify a reference process $\mathbb{Q}$ defined in Equation˜1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value. The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
@@
$m_{t}$ $=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right),\quad n_{t}=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}},$
$\sigma_{t}$ $=\eta_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}}$
where $\text{SNR}_{t},\alpha_{t},\sigma_{t}^{2}$ are all defined in the reference process $\mathbb{Q}$. In Equation˜3, $m_{t}$ and $n_{t}$ are coefficients that enforce the bridge’s dual endpoints: at $t=0,m_{0}=1,n_{0}=0\rightarrow z_{t}=z_{0}$; at $t=T,m_{T}=0,n_{T}=1\rightarrow z_{t}=z_{T}$.
@@ 4.1 Problem Formulation
We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ denotes the target dense signal map (e.g. depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image. In the following sections, we’ll elaborate the details of the DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
@@ 4.2 Dense Prediction Bridge
Given paired training data, a straightforward approach is to construct the diffusion bridge model from scratch, following the formulation in Sec.˜3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more, it cannot to harness rich visual prior embedded in the foundation models pretrained with large-scale data, which have been shown to significantly enhance both perceptual quality and generalization capability in dense prediction tasks. To address these limitations and fully exploit the visual prior, our DPBridge is built upon pretrained diffusion backbones through finetuning. We explore several strategies to handle the misalignment between bridge process and standard diffusion process, facilitating more efficient knowledge transfer. An overview of our training pipeline is provided in Figure˜2, followed by a detailed explanation of each component.
@@ 4.3 Latent Space Implementation
We follow common practice to perform diffusion steps in a compressed yet perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the dense signal map and RGB image are encoded using the VAE encoder, yielding latent representations $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, which serve as the initial and terminal points of the diffusion bridge process. In this way the diffusion bridge process is operated entirely in the latent space. To recover the original signal map, we use the VAE decoder to decode the predicted latent $\hat{\mathbf{y}}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. For clarity, all subsequent derivations in this paper are conducted in latent space.
@@ 4.4 Training Objective
Maximum Likelihood Loss. A widely adopted approach for learning the reverse process associated with Equation˜2 follows the theoretical framework of [anderson1982reverse], where a neural network is trained to approximate the score function [zhou2023denoising]. In contrast, we observe that in the diffusion bridge setting, the reverse transition kernel admits a tractable analytical form. This insight enables us to adopt a maximum likelihood training paradigm in place of score matching. Since widely used latent diffusion models like Stable Diffusion [rombach2022high] are pretrained under a likelihood-based loss function, finetuning with a consistent objective is more compatible, which helps preserve prior knowledge and improves training stability.
@@
$\emph{ELBO}=$ $\mathbb{E}_{p(\mathbf{z}_{0})}\Big[\mathbb{E}_{p(\mathbf{z}_{1}\mid\mathbf{z}_{0})}\big[\log p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{1},\mathbf{z}_{T})\big]$- (4)
@@
For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$. Based on Equation˜3, we further derive the mean value term:
@@
$k_{1}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t},\quad k_{2}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot m_{t-1},$
$k_{3}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
@@
Assuming $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ also follows a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ with fixed variance $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)=\tilde{\sigma}_{t}^{2}$ for simplicity of derivation. Following the strategy of DDPM [ho2020denoising], we reparameterize the model $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to predict noise instead of posterior mean. Under this formulation, the ELBO objective is equivalent to:
@@
While Equation˜8’s MSE-like form resembles DDPM’s noise loss, its derivation relies on the analytical reverse kernel of the bridge process. The inclusion of $z_{T}$ is not “standard conditional generation” but a requirement for enforcing the bridge’s dual-endpoint constraint. For more derivation details, please refer to Appendix˜A